OpenStudy (anonymous):

(secx + tanx)^2 = (1+sinx) / (1-sinx)

OpenStudy (anonymous):

wow, what a question, sstarica, you think you can handle it?

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

thinking :P

OpenStudy (anonymous):

i think you should use google first =D

OpenStudy (anonymous):

sstarica, if you become my fan, I will become yours, a win win

OpenStudy (anonymous):

\[= 1+sinx / 1-sinx\]\[= \sec^x + 2secxtanx + \tan^2x\]\[= 1 + \tan^2x + 2(sinx/\cos^2x) + 2(\sin^2x/\cos^2x)\]\[= 1 + ((2\sin^2x + 2sinx)/\cos^2x)\]\[= 1 + (2sinx(1+sinx)/1-\sin^2x)\]\[=1 + (2sinx(1+sinx)/(1-sinx)(1+sinx))\]\[=1 + (2sinx/1-sinx)\]\[= 1(1-sinx) + 2sinx/1-sinx\]\[= 1-sinx + 2sinx/1-sinx\] ^_^

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

= 1 + sinx / 1-sinx :)

OpenStudy (anonymous):

no need for google :)

OpenStudy (anonymous):

wow, you are smart

OpenStudy (anonymous):

2nd line should be sec^2 x and not sec^x ^^" typo sorry

OpenStudy (anonymous):

but I LOVE google =D

OpenStudy (anonymous):

we all are ^_^

OpenStudy (anonymous):

wait a min....ANGOO! :D

OpenStudy (anonymous):

that's you!

OpenStudy (anonymous):

Dr. GOO LOL!

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

another account? lol, I'm slow sorry ^^" just figure u out =P

OpenStudy (anonymous):

figured*

OpenStudy (anonymous):

why another account? bored? :D

OpenStudy (anonymous):

yes =D

OpenStudy (anonymous):

got a better name LOL

OpenStudy (anonymous):

I'll call you GOOBOO :D!!!!

OpenStudy (anonymous):

fanned you

OpenStudy (anonymous):

fine fine lol

OpenStudy (anonymous):

happy? =P

OpenStudy (anonymous):

yes, VERY =D

OpenStudy (anonymous):

seriously, lol

OpenStudy (anonymous):

ok, I got to go for now, book plane tickets for my dad

OpenStudy (anonymous):

see you later smart lady

OpenStudy (anonymous):

alright , see you then ^_^

OpenStudy (anonymous):

alright Dr.G